OpenStudy (anonymous):

\frac{ x-3 }{ 4 }=\frac{ 5 }{ 14 }-\frac{ x+5 }{ 7 }

OpenStudy (paxpolaris):

\[\frac{ x-3 }{ 4 }=\frac{ 5 }{ 14 }-\frac{ x+5 }{ 7 }\]

OpenStudy (paxpolaris):

multiply everything by 28

OpenStudy (paxpolaris):

\[7(x-3)=2(5)-4(x+5)\]

OpenStudy (anonymous):

then you distrubute